A leading well panel manufacturer had an year end promotion sale on last December. At the end of the sale they found out that the amount of sales they have done after "x" number of days was; show that the differential equation


1
Expert's answer
2021-04-13T16:12:52-0400

Since differential equation is not given in the question,

We can assume the differential equation, whose solution can be find out


"\dfrac{dx}{dt} = 0.02x"


"\int \dfrac{dx}{x} = \int 0.02dt"


"logx = 0.02t +C"


"x = e^{0.02t+C}"